OpenStudy (anonymous):

Find the domain and range of y = (squareroot) x + 4* Thanks

OpenStudy (anonymous):

domain if x + 4 is all under the radical then x+4 >=0 and solve for x Range look its graph in the calculator and see all the possible values y can be, Im pretty sure y>=4 but check on that. Dont have a graphing calculator? Buy one.

OpenStudy (anonymous):

if it is \[y=\sqrt{x+4}\] then the domain is \[x\geq-4\] and the range is \[y\geq 0\] because the radical symbol means positive square root